OpenStudy (anonymous):

Can someone put this into an equation? The length of a rectangle is 11 in. longer than its width. The perimeter of the rectangle is 86 in. What is the length of the rectangle? 38 in. 32 in. 27 in. 16 in.

OpenStudy (igreen):

Plug what you know in the Perimeter formula: \(P = 2L = 2W\)

OpenStudy (anonymous):

So what am I looking for?

OpenStudy (igreen):

\(86 = 2(11) + 2W\) Multiply: \(86 = 22 + 2W\) Subtract 22 to both sides: \(64 = 2W\) Divide 2 to both sides: \(W = 32\)

OpenStudy (anonymous):

OK, so the answer is B 32? I just got done with a class about this. This question just confused me. :)

OpenStudy (igreen):

Yes.